OpenStudy (anonymous):

Why is the exact value of tan^-1 (sqrt 3) and tan(sqrt 3) both equal to pi/3?

OpenStudy (zarkon):

they aren't

OpenStudy (anonymous):

but i always get the same exact value

OpenStudy (carson889):

Make sure you are using proper brackets and what not. On a site called wolframalpha.com, type in: tan^-1(sqrt(3)) and note the value. Then on the same site plug in: tan(sqrt(3)). Note and compare the values, they won't be the same. Also make sure you type the numbers in exactly as i have typed them.

OpenStudy (anonymous):

It didnt show the exact value for tan(sqrt(3)), so there is no exact value for it..?

OpenStudy (zarkon):

there is ...namely \[\tan(\sqrt{3})\] :)

OpenStudy (zarkon):

if you want a decimal approximation it is about -6.1475

OpenStudy (anonymous):

so...is tan (sqrt 3) = pi/3...?

OpenStudy (zarkon):

is \[\frac{\pi}{3}=-6.1475...\] ?

OpenStudy (anonymous):

oh..the CAST rule?

OpenStudy (anonymous):

but wouldn't tan^-1 be -ve as well.. omg..i really dont get it

OpenStudy (anonymous):

oh wait, nevermind . i think i get it now

OpenStudy (anonymous):

the arc tan is to use to find angle tan is for find value right

OpenStudy (zarkon):

\[\tan^{-1}(x)\] is the inverse tangent function we have \[\tan^{-1}(\sqrt{3})=\frac{\pi}{3}\] so \[\tan\left(\frac{\pi}{3}\right)=\sqrt{3}\]

OpenStudy (anonymous):

yea, i get it now thnks i always mix it up with 1/ tan

OpenStudy (anonymous):

so i thought u need to flip to oppo / adjacent

OpenStudy (anonymous):

from opp/adja to adja/ opp now i am clear, thx

OpenStudy (zarkon):

good